OpenStudy (anonymous):

The standard form of a parabola is y=ax^2+bx+c , but what I don't get is what each letter represents.

OpenStudy (anonymous):

'x' is the variable. It's value you input to get an output value 'y'. For any parabola, the domain of 'x', meaning the x-values that you are allowed to input are said to be all REAL NUMBERS. Now 'a' and 'b' are constants and they are the coefficients of x^2 and x respectively. 'c' is just a constant that hangs around at the end. Here is an example equation of a parabola:\[\bf y=2x^2+3x-6\]Here a = 2, b = 3, c = -6.

OpenStudy (anonymous):

If I wrote:\[\bf y=x^2\]Then a = 1, b = 0, and c = 0. You can see this by re-writing the x^2 as:\[\bf y=x^2+bx+c\]Since we know that the equation is just "y = x^2" then the bx and c must cancel. This will only happen if b = c = 0.

OpenStudy (anonymous):

Okay, thank you so much! It's coming back to me now! :D